My son is in pre-algebra. I am trying to help him, but I don't know what this is called! I did the questions in excel and came out with the right answers, but I don't know how to get there! I am exacerbated! Anyway...I know I am missing something, but what? -3s+2(-5s+1) s=-3
Ok, you solve what is in the ( ) first (right?); so (-5s+1) is (-5*-3+1)=-5*-3=15 then 15+1=16; so -3s+2(16); then -3*-3=9; 9+2(16), 11(16), 11*16=the wrong answer! How do you get 41? Please help. It has been over 20 years since I did any of this, and I don't remember any of it!
Thank you,
Shelli

Answer by kalants(7)

Well, you've got all the right steps 'til the end. You're multiplying instead of adding.
-3s+2(-5s+1) s=-3
-3(-3)+2(-5(-3)+1)=
9+2(15+1)=
9+2(16)=
9+32=
41
Hope that helps...